### v1

# CoTAR: Chain-of-Thought Attribution Reasoning with Multi-level Granularity

Abstract
State-of-the-art performance in QA tasks is currently achieved by systems employing Large Language Models (LLMs), however these models tend to hallucinate information in their responses. One approach focuses on enhancing the generation process by incorporating attribution from the given input to the output. However, the challenge of identifying appropriate attributions and verifying their accuracy against a source is a complex task that requires significant improvements in assessing such systems. We introduce an attribution-oriented Chain-of-Thought reasoning method to enhance the accuracy of attributions. This approach focuses the reasoning process on generating an attribution-centric output. Evaluations on two context-enhanced question-answering datasets using GPT-4 demonstrate improved accuracy and correctness of attributions. In addition, the combination of our method with finetuning enhances the response and attribution accuracy of two smaller LLMs, showing their potential to outperform GPT-4 in some cases.
CoTAR: Chain-of-Thought Attribution Reasoning with Multi-level Granularity
Moshe Berchansky   Daniel Fleischer   Moshe Wasserblat   Peter Izsak Intel Labs {moshe.berchansky,daniel.fleischer,moshe.wasserblat,peter.izsak}@intel.com
1 Introduction
Text generation from sources, such as information obtained through retrieval, is a key aspect of grounded question answering. One way to ensure the credibility of these models is through attributed text generation. This method pairs the generated text with supporting evidence, enhancing the model’s trustworthiness and allowing for easier detection of errors.
The pairing process can be done in several levels. Firstly, we can pair each sentence in our response with a set of passage identifiers Gao et al. (2023). On a model detailed level, individual spans in the answer can correspond to a specific passage, hence attributing specific portions to various passages Schuster et al. (2023). These spans are designed to be direct copies from the passages, assuring that the content provided is fully supported. Since not all the spans are copied, the model has a sufficient degree of flexibility to produce both a coherent and factually correct answer.
Figure 1: Usage of CoT for attribution-based answers.
However, these approaches fail at times in two primary ways to cite accurately. On one hand, the model might focus only a very specific portion of the input, therefore missing relevant sections when it constructs the answer. On the other hand, it might cite too many passages, some of which might not be relevant to the answer.
To solve these, we propose utilizing a Chain-of-Thought approach (CoT) Wei et al. (2022b), which we denote as CoTAR, which allows the model to perform reasoning over the input passages before generating the output. In our approach, we instruct the model to extract relevant information from the passages as well as specify the form of attribution to generate. The information can be in the span, sentence, or passage level. We inspect how the various CoT methods affect the model’s output, in each of the citation levels. In addition to few-shot instruction, we also finetune various models using our method, resulting in models that are competitive with and even outperform GPT-4 in some cases.
We can summarize our contributions in this paper as follows:
  • We perform rigorous measurements of both the answer quality and citation quality, across multiple models and citation levels.
  • We show that utilizing CoT reasoning improves the ability of an LLM to produce both better quality answers, and more precise and faithful citations from the source, demonstrated on multiple models.
  • We demonstrate that using by finetuning, smaller models can be competitive with or outperform GPT-4 in some cases in answer and citation quality metrics.
2 Method
2.1 Attribution-Oriented Question Answering
Attribution-oriented question answering can be defined as a task where, given a query and a set of relevant contextual resources, the objective is to accurately answer the question while attributing specific portions or the entire answer to the appropriate contextual sources.
We can identify three levels of attribution with different granularity levels: Span, Sentence and Passage. Attribution levels refer to how citations are displayed within an answer, indicating whether complete passages, individual sentences, or specific text segments are credited back to their original sources. Detailed examples for each level are available in Table 4 in the appendix.
2.2 Chain-of-Thought Attribution Reasoning
We propose a multi-step CoT Wei et al. (2022b) reasoning scheme with varying levels of attribution, similar to the three levels presented in Section 2, hypothesizing that this could encourage the model to generate more accurate answers. The process involves identifying the most crucial aspects of the given context for answering the question, by incorporating direct citations to the referenced parts. We denote the process of granular attribution as CoT attribution guidance, or CoT method for short. The three levels are as follows:
  • Span Guidance: Produce the relevant spans of information per passage.
  • Sentence Guidance: For every passage, write sentences that summarize how the passage answers the question.
  • Passage Guidance: State which passages are relevant for the question.
We explore all combinations of CoT methods with citation levels, and asses how each one effects the generated answers. A example of all three levels of attribution-guidance is in Table 5 in the appendix.
Level CoT Answer Quality Citation Quality BERT HEM RL Avg. ALCE F1 CSCA DOC F1 SEM-F1 Avg. Span None 91.71 67.58 66.08 75.12 77.06 81.15 85.33 67.57 77.78 Span 92.06 66.95 68.55 75.85 78.83 89.88 86.54 72.23 81.87 Sent. 91.87 66.88 68.76 75.84 84.76 69.69 86.38 70.52 77.84 Pass. 91.91 66.89 67.67 75.49 79.21 86.94 86.49 69.87 80.63 Sent. None 91.47 62.48 62.18 72.04 82.79 - 85.32 59.20 75.77 Span 91.92 62.06 64.76 72.91 82.39 - 86.71 61.74 76.94 Sent. 91.96 64.59 67.36 74.64 86.42 - 86.51 64.44 79.12 Pass. 91.69 61.19 64.42 72.43 84.25 - 85.92 61.79 77.32 Pass. None 91.55 61.13 62.76 71.81 75.38 - 85.19 44.31 68.29 Span 91.90 61.66 64.70 72.75 71.60 - 86.17 43.41 67.06 Sent. 91.84 62.43 65.32 73.20 70.63 - 86.41 42.59 66.54 Pass. 91.69 60.40 63.88 71.99 73.28 - 85.64 44.12 67.68
Table 1: Results of GPT-4 on the QuoteSUM dataset. The results are categorized by response attribution (level), and for each level, every CoT approach is demonstrated. The best results per level are marked in bold.
3 Evaluation Metrics
In order to evaluate the answers, we specify metrics of two types. Answer Quality, measuring similarity between the predicted answer and the gold answer. Citation Quality, assessing the similarity between the cited text, the cited passage, and the citations present in the gold answer.
For answer quality , we use the n-grams based method ROUGE-L (RL), and the semantic method BERTScore (BERT) Zhang* et al. (2020). For hallucination evaluation, we use the HEM111vectara/hallucination_evaluation_model model, which was finetuned on NLI datasets, and predict whether two texts are factually consistent. For citation quality, we measure the quality of the cited content; we propose specific metrics for each of the citation levels. We note that span citations must match the source passages precisely, while sentence and passage attributions are not required to do so. We use the SEM-F1 metric proposed by Schuster et al. (2023) for n-gram token level similarity between the cited content. In addition, we include the Citation Precision/Recall introduced in Gao et al. (2023), and combine them into ALCE F1, by using the harmonic mean. We also measure the F1 score over passage indices between the cited passages and the passages cited by the expected answer, denoted as DOC F1, and the Correct Span Citation Attribution (CSCA) which indicates whether a predicted span is a direct span from the attributed passage.
4 Experimental Setup
4.1 Datasets
We use the QuoteSum (QSUM) Schuster et al. (2023) dataset, which contains semi-extractive answers written by humans, for natural questions and generated ones. In addition, we include MS MARCO (MS) [^2] https://huggingface.co/datasets/ms_marco Campos et al. (2016), which is a crowd sourced dataset of responses to queries from Bing. Each example include 10 passages which can be relevant or irrelevant to answering the question.
4.2 Span Dataset Preparation
In order to include span citations in datasets such as MSMARCO, we can create span-based citation labels out of passage-based labels. We do so by finding the longest common substrings between the input passages and the answer, and choosing those that contain named entities, thus ensuring the semantic relevance of the cited spans. The answer is then formatted as proposed by Schuster et al. (2023). For more details about the process, we refer the reader to Algorithm 1.
4.3 Models
We use GPT-4 Achiam et al. (2023) for evaluating our CoT method. We use the Mistral 7B Jiang et al. (2023) decoder model and the Flan-T5 XXL Chung et al. (2022) encoder-decoder as representative small-sized models in our further finetuning evaluations.
4.4 Method Comparison
For each of the tasks described in Section 2.1, we utilize every CoT method described in Section 2.2. We run each combination with GPT-4, a small decoder-only model, and an encoder-decoder model. All results are presented on the test sets of the datasets mentioned in Section 4.1. For full implementation details of the training and inference, we refer the reader to section A in the Appendix.
5 Experimental Results
To examine how CoT affects the model, we inspect the GPT-4 case, and provide a detailed overview of the metrics in Table 1. In the answer metrics, the differences between the various CoT methods and the standard run (shown as "None") are not significant, aside from the sentence level, where the sentence CoT method produces better scores across all metrics. For the citation quality metrics, the span and sentence level produce better scores than their counterparts when using the appropriate CoT methods. In particular, the CSCA metric in the span level case showcases a significant advantage, mainly since the model is guided to produce the spans relevant for the answer in advance, thus allowing the model to accurately reference them. At the passage level, the SEM-F1 and ALCE-F1 scores showcase an advantage for the standard run. We hypothesize that the addition of CoT for such a broad citation task is less helpful than in the more granular sentence and span level citations. We observe that the GPT-4 model performed best in the span level when using the span CoT method, and the same applies for the sentence level with sentence CoT. In the passage level, while the CoT methods improve the answer quality, they do not aid in the citation quality.
We showcase similar results for all the models, across the various citation levels and CoT methods in Tables 7 and 8. For each setting, we display all metrics and category averages for answer and citation quality, respectively.
We present an overview of the various combinations of citation levels and CoT methods in Table 2. For smaller decoder-only models (i.e. Mistral 7B), the answer scores perform best in general when the citation level is used with its appropriate CoT method. In the citation metrics, results are different between the QSUM and MS. While the CoT methods behave well in the MS dataset, they are proving less helpful in QSUM, with the main difference being a higher ALCE-F1 score for the standard run. The difference can be mostly attributed to the higher complexity of the QSUM dataset, as opposed to the shorter and simpler MS answers.
For encoder-decoder models, the CoT methods produce better results than the standard run, both in the answer and citation metrics. However, aside from the MS citation quality results, the advantage of using each citation level task with its corresponding CoT method is less prominent. The sentence CoT method is highly dominant in most of the metrics in the QSUM dataset, and passage CoT method being prominent in the MS dataset.
Finally, we observe in Table 2 that the smaller models are competitive with and even outperform GPT-4. This conclusion is similar to the one reached by Schuster et al. (2023), when comparing to other LLMs without finetuning.
CoT Answer Quality Citation Quality
Span Sent. Pass. Span Sent. Pass.
QuoteSUM
GPT-4 None 75.1 72.0 71.8 77.8 75.8 68.3
Span 75.9 72.9 72.8 81.9 76.9 67.1
Sent. 75.8 74.6 73.2 77.8 79.1 66.5
Pass. 75.5 72.4 72.0 80.6 77.3 67.7
Mistral 7B None 74.2 72.6 74.4 83.0 72.7 66.4
Span 76.2 75.6 75.1 82.2 72.3 64.7
Sent. 75.2 75.6 75.1 79.9 73.4 65.0
Pass. 75.6 74.8 75.8 82.0 73.5 66.1
Flan-T5 None 76.6 73.8 74.0 84.7 71.4 64.3
Span 77.3 71.8 70.6 84.1 69.7 62.1
Sent. 78.4 78.2 76.7 84.7 79.7 66.5
Pass. 78.5 74.4 73.1 85.6 72.7 61.2
MSMARCO
Mistral 7B None 69.6 69.1 70.2 68.4 66.1 66.0
Span 70.8 70.5 71.5 74.8 67.8 68.9
Sent. 71.0 71.6 71.3 72.4 69.1 68.4
Pass. 70.8 71.6 71.6 74.4 69.2 69.1
Flan-T5 None 70.4 71.5 72.4 74.3 67.7 68.5
Span 70.4 70.3 68.4 74.7 68.7 68.6
Sent. 71.3 71.6 71.9 73.2 69.0 69.3
Pass. 69.0 72.7 72.6 73.9 69.0 70.0
Table 2: Average answer and citation quality results, per citation level, per CoT method, for QuoteSUM and MSMARCO. Rows: Citation levels. Columns: CoT method used and metrics averages per level.
6 Related Work
Several recent studies have suggested different techniques for text generation including citation attribution, with varying degrees of detail. Attributing relevant information to web pages was done in Thoppilan et al. (2022); Liu et al. (2023). Menick et al. (2022) attributes the response to specific snippets from the source passages. In particular, Gao et al. (2023) assigns attribution for each produced sentence to one or several input passages. Our sentence level citation approach and the evaluation metrics are based on this work.
Substantial work has been done to reduce hallucination in the generated text produced by the model Li et al. (2024); Rawte et al. (2023), as citation has the ability to increase the reliability of the generated text. Some work has been done on correlating answer to text sections in the retrieved passages Gao et al. (2023); Bohnet et al. (2023). However, these techniques require additional processing.
The usage of CoT has shown significant improvements in a wide variety of tasks Wei et al. (2022b). Additionally, in-context learning can be added to the input, by providing examples of how to solve the task at hand Wei et al. (2022a). We explore the effects of using CoT with the various citation level tasks, in addition to the usage of in-context learning examples in every scenario. Luo et al. (2023) instruct and finetune models to produce claims, based on retrieved passages, differentiating between informative and distracting input segments. We specify the text granularity level we expect the model should cite and phrase its answers. Slobodkin et al. (2024) focus on the sentence citation task, splitting up the generation task into multiple reasoning steps, including CoT as well. Our approach focuses on using a single call, producing the answer at-once, with a CoT approach.
7 Conclusions
Through extensive measurements applied across various models and citation levels, we have been able to ascertain the quality of both answers and citations. Our findings indicate that the use of our CoTAR reasoning significantly enhances the capacity of a model to generate superior quality answers and more accurate, faithful citations from the source. This improvement is evident in both decoder-only and encoder-decoder models. Furthermore, our research shows that by using finetuning, smaller models can compete with or even surpass the performance of GPT-4 across a wide range of answer and citation metrics. This is particularly evident when considering the diversity of citation levels and CoT methods.
8 Limitations
In our experimentation, we did not compare our approach directly to the work done in Slobodkin et al. (2024), due to lack of access to the publicly available code-base. In our GPT-4 runs, we ran the experimentation solely on the QuoteSUM dataset, in order to compare GPT-4 to our models on an previously established dataset. When considering datasets for our work, we primarily focused on the QuoteSUM dataset from Schuster et al. (2023), and also the MSMARCO dataset, due to the extensive annotation of the answers and the accompanying passages. For future work, we will add similar datasets in structure and content.
9 Potential Risks
Since this project is primarily focused on assessing current models using datasets that are openly accessible, we do not foresee any possible adverse effects.
Appendix A Implementation Details
The prompt is built using 4 fewshot examples, each composed of a question, a list of passages, and the expected answer in the task’s format. As done by Schuster et al. (2023), the prompt for the models is the same as they have specified, with an example in Table 6, and the fewshot examples are chosen with the question similarity between the test question and the training example questions. As stated above, the expected answers cite their passages of origin in one of the levels described in Table 4. For the CoT setting, a prefix is attached in one of the methods showcased in Table 5. Text generation is done without sampling, with a maximum length of 2k tokens. For QuoteSUM, the train and test sets are identical to the ones used by Schuster et al. (2023). For MSMARCO, we utilize 2750 samples from the created dataset for training, and a 1000 for the test set. We specify the training parameters in Table 3.
Appendix B Models
The weights for the models we have used have all been retrieved from HuggingFace, aside from GPT-4, for which we use Azure OpenAI:
Appendix C Additional Results
Parameter Decoder-Only Encoder-Decoder
Max Target Len. - 4096
Max Seq. Len. 8192 32768
Lora R 16 16
Lora $\alpha$ 32 32
Lora Dropout 0.05 0.05
Lora Bias None None
Lora Modules gate, down, up, $q$, $v$, $k$, $o$ $q,v$
LR 5e-5 2e-4
LR Scheduler Linear Linear
Weight Decay 0 0.01
Precision bfloat16 bfloat16
Batch Size 1 1
Epochs 3 3
Warmup Ratio 0 0.1
Table 3: The training parameters used for Lora finetuning for the decoder-only and encoder-decoder models.
Algorithm 1 Generating span attirubtions from a passage-labeled dataset.
1:Input: Answer $a$, passages $\{p_{i}\}$
2:Extract named entities $\text{NER}(a)$ using spaCy Honnibal and Montani (2017).
3:Find common substrings $CS(a,p_{i})$ between every passage $p_{i}$ and answer $a$.
4:Each substring $sc\in CS(a,p_{i})$: contains $sc_{t}$ (string text) and passage index $sc_{p}$.
5:Sort substrings by length across all passages into $SC(a)$ in a descending order.
6:$a_{\text{marked}}$: A dictionary that maps the the letters of the answer $a$ to either "marked" or "unmarked".
7:$\text{IsMarked}(sc,a,a_{\text{marked}})$: Returns true if the string $sc$ has already been marked or marked partially inside answer $a$, and false otherwise.
8:$\text{Mark}(sc,a,a_{\text{marked}})$: Marks the string in the answer as "marked" under passage $sc_{p}$.
9:for $sc\in SC(a)$ do
10:     if  $!\text{IsMarked}(sc,a,a_{\text{marked}})$ then
11:         $\text{Mark}(sc,a,a_{\text{marked}})$
12:     end if
13:end for
14:Format marked spans stored in $a_{\text{marked}}$ with the answer $a$ to form $a_{\text{span}}$.
15:return $a_{\text{span}}$
Citation Level Cited Answer
Passage
" Johnny Panic and the Bible of Dreams " is a song by the British
band Tears for Fears. They also have international hit singles
such as: "Mothers Talk", "Shout", "Everybody Wants to Rule the World",
"Head over Heels", and "I Believe" [1][5].
Sentence
" Johnny Panic and the Bible of Dreams " is a song by the British
band Tears for Fears [1]. They also have international hit singles
such as: "Mothers Talk", "Shout", "Everybody Wants to Rule the World",
"Head over Heels", and "I Believe" [5].
Span
" [ 1 Johnny Panic and the Bible of Dreams ] " [ 1 is a song by the British
band Tears for Fears ] . They also have [ 5 international hit singles ]
such as: [ 5 "Mothers Talk", "Shout", "Everybody Wants to Rule the World",
"Head over Heels", and "I Believe". ]
Table 4: Citation Level examples, with each row highlighting the approach for citing the source passages.
CoT Method Answer Prefix
Passage
Lets analyze the input passages.
The only relevant passages to the question are passages 1, 5.
Thus, the final answer is:
Sentence
Lets analyze the relevant information from the input passages.
From passage [1], we know that: " Johnny Panic and the Bible of Dreams "
is a song by the British band Tears for Fears .
From passage [5], we know that: They also have international hit singles
such as: "Mothers Talk", "Shout", "Everybody Wants to Rule the World", "Head
over Heels", and "I Believe".
Thus, the final answer is:
Span
Lets analyze the relevant spans of information from the input passages.
From passage [1], the relevant spans are the following:
Johnny Panic and the Bible of Dreams
is a song by the British band Tears for Fears
From passage [5], the relevant spans are the following:
international hit singles
"Mothers Talk", "Shout", "Everybody Wants to Rule the World", "Head
over Heels", and "I Believe".
Thus, the final answer is:
Table 5: Examples of Chain-of-Thought prompt prefix for the various Levels. The model is instructed/finetuned to produce the snippets per CoT method.
Answer the question by summarizing the given sources while explicitly copying
spans from the sources. When copying a span, use brackets and the respective
source number to indicate that this span was copied. Use explicit copying
as much as possible and for all factual statements, while preserving fluency.
Make sure to use all relevant sources and properly quote them. Here
are some examples:
Question: how much power does a wind turbine produce?
[1] Compact wind acceleration turbine: It is generally thought that since…
[2] Sustainable architecture: roof ledge. Small-scale rooftop wind turbines have …
[3] Turby wind turbine: can because horizontal axis (HAWT) types cannot change …
Quoted summary: One source states the [ 1 amount of power produced by a wind
turbine is proportional to the cube of the wind speed ] . Other sources state
[ 2 Turbines for residential scale use ] [ 2 produce electricity at a rate of 900 watts to
10,000 watts ] , and [ 3 is specified to generate power in winds of
between 4 m/s (9 mph, 7.8kts) and 14 m/s (31 mph, 27.2kts) ] .
Question: a component is what?
[1] Modular programming: in Dart, Go or Java) is sometimes used instead of …
[2] Physical body: the system at a point in time changes from identifying the object to…
Quoted summary: A [ 1 component is a piece of a whole system ] . Also, [ 2 A
component is an object completely within the boundary of a containing object. ]
Question: what is the title of tears for fears song?
[1] Johnny Panic and the Bible of Dreams (song): …
[2] Mark Crew: Will Ruin Your Life. The album was produced at his studio. …
[3] Raoul and the Kings of Spain: Raoul and the Kings of Spain is the fifth studio …
[4] Everybody Loves a Happy Ending: Everybody Loves a Happy Ending …
[5] Songs from the Big Chair: Songs from the Big Chair is the second studio album …
Quoted summary: Song titles by [ 1 the British band Tears for Fears ] include the
[ 5 1985 ] [ 5 international hit singles "Mothers Talk", "Shout", "Everybody Wants to
Rule the World", "Head over Heels", and "I Believe" ] , [ 4 1989’s "The Seeds of Love" ]
, the [ 1 1990 single "Advice for the Young at Heart" ] with [ 1 "Johnny Panic and the Bible
of Dreams" ] on [ 1 the B-side ] , and " [ 2 Lemon To A Knife Fight ] ",
" [ 2 Cheetah Tongue ] ", and " [ 2 Turn ] " from [ 2 2018 ]
Table 6: Citation Level prompt example, containing 2 fewshot examples, with the relevant passages for each question provided.
Level CoT Model Answer Quality Citation Quality
BERT HEM RL AVG. ALCE F1 CSCA DOC F1 SEM-F1 AVG.
Span GPT-4 91.71 67.58 66.08 75.12 77.06 81.15 85.33 67.57 77.78
None Flan 92.20 65.41 72.12 76.58 85.92 94.26 86.71 71.94 84.71
Mist. 91.61 60.93 70.03 74.19 86.95 94.97 84.69 65.27 82.97
GPT-4 92.06 66.95 68.55 75.85 78.83 89.88 86.54 72.23 81.87
Span Flan 92.06 67.97 71.99 77.34 81.47 95.36 86.51 72.86 84.05
Mist. 92.04 65.98 70.67 76.23 74.60 96.17 85.87 72.19 82.21
GPT-4 91.87 66.88 68.76 75.84 84.76 69.69 86.38 70.52 77.84
Sent. Flan 92.15 71.46 71.64 78.42 86.71 90.02 87.79 74.41 84.73
Mist. 92.01 63.32 70.22 75.19 73.56 90.86 85.40 69.69 79.88
GPT-4 91.91 66.89 67.67 75.49 79.21 86.94 86.49 69.87 80.63
Pass. Flan 92.17 70.84 72.47 78.49 83.36 94.14 88.01 77.07 85.64
Mist. 91.88 63.90 70.91 75.56 80.14 94.12 85.03 68.63 81.98
Sent. GPT-4 91.47 62.48 62.18 72.04 82.79 - 85.32 59.20 75.77
None Flan 91.84 60.43 69.22 73.83 79.37 - 81.18 53.66 71.40
Mist. 91.56 58.51 67.63 72.57 82.15 - 82.10 54.00 72.75
GPT-4 91.92 62.06 64.76 72.91 82.39 - 86.71 61.74 76.94
Span Flan 91.43 57.43 66.60 71.82 71.33 - 83.19 54.46 69.66
Mist. 91.93 65.04 69.69 75.55 73.14 - 84.48 59.29 72.30
GPT-4 91.96 64.59 67.36 74.64 86.42 - 86.51 64.44 79.12
Sent. Flan 92.37 69.36 72.87 78.20 85.38 - 87.03 66.77 79.72
Mist. 91.94 64.45 70.40 75.60 75.48 - 84.81 59.97 73.42
GPT-4 91.69 61.19 64.42 72.43 84.25 - 85.92 61.79 77.32
Pass. Flan 91.64 62.97 68.62 74.41 72.57 - 85.05 60.43 72.68
Mist. 91.99 62.90 69.37 74.76 75.39 - 85.18 60.07 73.55
Pass. GPT-4 91.55 61.13 62.76 71.81 75.38 - 85.19 44.31 68.29
None Flan 91.69 62.10 68.26 74.02 69.33 - 83.31 40.41 64.35
Mist. 91.74 63.64 67.91 74.43 75.30 - 83.93 40.08 66.44
GPT-4 91.90 61.66 64.70 72.75 71.60 - 86.17 43.41 67.06
Span Flan 91.24 54.87 65.56 70.56 65.51 - 81.98 38.76 62.09
Mist. 91.99 64.25 69.14 75.12 70.65 - 83.52 39.78 64.65
GPT-4 91.84 62.43 65.32 73.20 70.63 - 86.41 42.59 66.54
Sent. Flan 92.43 65.17 72.61 76.73 73.00 - 84.47 42.05 66.51
Mist. 92.14 62.54 70.59 75.09 69.92 - 84.53 40.61 65.02
GPT-4 91.69 60.40 63.88 71.99 73.28 - 85.64 44.12 67.68
Pass. Flan 91.32 61.82 66.14 73.09 58.18 - 86.16 39.20 61.18
Mist. 92.15 64.54 70.83 75.84 72.92 - 84.64 40.76 66.11
Table 7: Results for the GPT-4 with fewshot and the smaller fintuned models over the QuoteSUM dataset. Each row contains both the values of the metrics per citation level and CoT method, and the average score on per metric type on their right. Mist. denotes Misrtal-7B, and Flan denotes FlanT5-XXL.
Level CoT Model Answer Quality Citation Quality
BERT HEM RL AVG. ALCE F1 CSCA DOC F1 SEM-F1 AVG.
Span None Flan 91.91 62.17 57.08 70.39 96.97 91.88 64.70 43.80 74.34
Mist. 91.24 64.16 53.28 69.56 80.61 94.51 59.86 38.47 68.36
Span Flan 91.80 62.83 56.50 70.38 95.69 93.58 64.90 44.72 74.72
Mist. 91.88 63.77 56.63 70.76 93.41 96.47 65.30 43.84 74.75
Sent. Flan 92.25 62.72 58.80 71.26 95.67 86.94 65.50 44.75 73.22
Mist. 91.83 64.31 56.85 70.99 93.38 91.62 63.40 41.18 72.40
Pass. Flan 91.90 56.24 58.77 68.97 95.93 90.09 64.90 44.55 73.87
Mist. 91.90 63.69 56.73 70.77 93.73 95.36 65.50 43.05 74.41
Sent. None Flan 92.83 62.77 58.78 71.46 96.53 - 62.00 44.49 67.68
Mist. 92.06 60.26 54.90 69.08 91.67 - 63.62 42.99 66.09
Span Flan 92.37 62.20 56.42 70.33 95.92 - 65.30 44.74 68.65
Mist. 92.54 61.66 57.24 70.48 94.75 - 64.50 44.09 67.78
Sent. Flan 92.81 63.31 58.82 71.65 96.55 - 65.15 45.29 69.00
Mist. 92.59 64.11 58.04 71.58 96.95 - 65.45 45.01 69.14
Pass. Flan 92.92 65.69 59.45 72.69 96.73 - 64.70 45.59 69.01
Mist. 92.67 63.88 58.20 71.58 96.05 - 66.40 45.14 69.19
Pass. None Flan 92.96 64.38 59.93 72.42 95.20 - 64.30 45.94 68.48
Mist. 92.29 62.06 56.32 70.22 92.54 - 62.80 42.75 66.03
Span Flan 92.17 55.69 57.29 68.38 96.10 - 65.00 44.59 68.56
Mist. 92.67 63.72 58.04 71.48 95.40 - 66.40 44.77 68.86
Sent. Flan 92.68 64.91 58.02 71.87 96.70 - 65.80 45.44 69.31
Mist. 92.49 64.12 57.16 71.26 96.30 - 64.80 44.20 68.43
Pass. Flan 92.97 65.20 59.49 72.55 96.60 - 67.40 46.10 70.03
Mist. 92.75 64.03 58.14 71.64 96.80 - 65.60 44.78 69.06
Table 8: Results for the smaller finetuned models over the MSMARCO dataset. Each row contains both the values of the metrics per citation level and CoT method, and the average score on per metric type on their right. Mist. denotes Misrtal-7B, and Flan denotes FlanT5-XXL.